Suppose the country of Macecon produces only two final goods, Theorems (Ts) and Lemmas (Ls). In 2018 it produced 10 Ts at a price of $50 and 50 Ls at a price of $10. In 2019 it produced 9 Ts at a price of $60 and 60 Ls at a price of $11. Using 2018 as the base year, the inflation rate in the GDP deflator between 2018 and 2019 was 5%.
1
Expert's answer
2020-04-16T09:52:02-0400

NominalGDP2019"=9\u00d760+60\u00d711=" $ "1200;"

RealGDP2019"=9\u00d750+60\u00d710=" $"1050;"

GDP deflator 2019"=\frac {1200}{1050}×100=" $ "114;"

Answer: GDP deflator is 14%.